# Template Matters: Understanding the Role of Instruction Templates in Multimodal Language Model Evaluation and Training

### v1

@@ Abstract
Current multimodal language models (MLMs) evaluation and training approaches overlook the influence of instruction format, presenting an elephant-in-the-room problem. Previous research deals with this problem by manually crafting instructions, failing to yield significant insights due to limitations in diversity and scalability. In this work, we propose a programmatic instruction template generator capable of producing over 39B unique template combinations by filling randomly sampled positional synonyms into weighted sampled meta templates, enabling us to comprehensively examine the MLM’s performance across diverse instruction templates. Our experiments across eight common MLMs on five benchmark datasets reveal that MLMs have high template sensitivities with at most 29% performance gaps between different templates. We further augment the instruction tuning dataset of LLaVA-1.5 with our template generator and perform instruction tuning on LLaVA-1.5-7B and LLaVA-1.5-13B. Models tuned on our augmented dataset achieve the best overall performance when compared with the same scale MLMs tuned on at most 75 times the scale of our augmented dataset, highlighting the importance of instruction templates in MLM training. The code is available at https://github.com/shijian2001/TemplateMatters.
@@ 1 Introduction
Multimodal Language Models (MLMs) have revolutionized vision-language learning by performing visual instruction tuning on diverse, high-quality multimodal instruction data [30, 64, 17, 21, 61]. MLMs achieve unprecedented performance on various visual tasks [26, 37, 39, 58, 55]. However, previous MLM evaluation and training methods overlook a significant elephant-in-the-room problem: different instruction formats will largely influence MLMs’ performance. Although recent studies [60, 56, 32, 46] demonstrate that MLMs may produce distinct outputs when changing the instruction format (as shown in Figure 1), research on MLMs’ sensitivity to instruction formats remains largely unexplored. Previous works designed hand-crafted instructions in a limited amount, which restricts the evaluation scale, thereby weakening their conclusions, and limiting the opportunity to finetune the MLMs with augmentation on different instruction formats.
Figure 1: An example of using different instruction templates to prompt MLM without changing the original QA pairs. The instruction templates are marked in blue. Prompting MLM with different instruction templates can twist the output of MLM.
@@
To systematically investigate the instruction sensitivity of MLMs and their impact on faithful evaluation, we propose to evaluate MLMs on Visual Question Answering (VQA) data augmented by various instruction templates without changing the meaning of the original QA pairs. To efficiently create diverse, high-quality instruction templates in sufficient quantities, we introduce a programmatic template generator that leverages diverse meta templates to produce semantically equivalent instruction templates automatically and scalably. Our approach can construct diverse instruction templates by random sampling from carefully curated word and phrase spaces to populate predefined placeholders, enabling the efficient generation of semantically consistent yet diverse instruction templates at scale. Our method can produce an extensive template space comprising 15K visual question templates and 249K choice-related templates, culminating in a comprehensive VQA instruction template space of 39B unique combinations. To effectively manage this vast template space, we use a tree-based organizational framework based on grammatical structures complemented by an efficient diversity sampling algorithm. This programmatic approach ensures the generation of instruction templates that maximize diversity across multiple dimensions, including grammatical construction, lexical choice, and symbolic representation.
We conduct a comprehensive robustness evaluation of instruction templates with our programmatic template generator, encompassing eight commonly used MLMs. Our experiment results reveal that those MLMs are highly sensitive to instruction template perturbation, with at most 29% performance gap across 100 different templates. We present the performance gap across instruction templates on the MMBench dataset in Figure LABEL:fig:banner(a). Given these results, we further introduce a simple yet effective method to improve visual instruction tuning that leverages our template generator to augment instruction datasets. We finetune two common MLMs (LLaVA-1.5-7B and LLaVA-1.5-13B) [28] using our generated diverse instruction templates and compare them with other MLMs finetuned on a larger scale (at most 75.19x than ours) of instruction-tuning datasets. Our finetuned MLMs achieve the best overall performance, demonstrating our method’s capability to improve MLMs in a data-efficient and low-cost way. We show the comparison of our 7B model to other models of similar scale on the SeedBench dataset in Figure LABEL:fig:banner(b). Our analysis further shows that compared to the original model, after finetuning with our template augmented instruction data, the model’s variance drops significantly on various out-of-domain instruction templates, which are not included in the training. Our approach not only validates the practical utility of our template generation framework but also illuminates promising directions for efficiently improving MLMs. On the other hand, our ablation studies show that models achieve the best general capabilities at a specific ratio between templates and training data, which varies with the model scale. We summarize our main contributions as follows.
@@
  • We introduce a novel programmatic instruction template generator that enables fast and scalable generation of diverse, semantically equivalent instruction templates.
  • We evaluate the robustness of eight commonly used MLMs to instruction format variations across five benchmarks leveraging our template generator, revealing their high sensitivity to instruction format variations.
@@
  • We propose a simple yet effective approach to enhance visual instruction tuning by augmenting the origin instruction-tuning dataset with programmatically scaling instruction templates. Our extensive experiments demonstrate its effectiveness.
@@ 2 Programmatically Scaling Instruction Templates
In this work, we propose a programmatic instruction template generator that efficiently produces a diverse array of grammatically correct and semantically consistent instruction templates without modifying the original Question-Answer (QA) pairs. Specifically, we generate instruction templates by programmatically populating placeholders in diverse meta templates with randomly sampled positional synonyms (phrases) to ensure flexibility while keeping the original meaning (Sec. 2.1). We organize our meta templates in a sentence pattern tree (Sec. 2.2), along with weighted sampling to ensure the sampling probability across all meta templates is uniformly distributed.
@@ 2.1 Meta Templates
A meta template $p_{i},i\in\{1,...,N\}$, serves as a formal blueprint for constructing instruction templates, consisting of a sequence of fixed string segments interspersed with placeholder $\langle h^{(i)}_{j}\rangle,j\in\{1,...,M_{i}\}$, where $M_{i}$ is the number of placeholders. We associate each placeholder $\langle h^{(i)}_{j}\rangle$ with a predefined set of synonyms (phrases) $s_{j}^{(i)}$. We design $s_{j}^{(i)}$ according to the semantic position of $\langle h^{(i)}_{j}\rangle$, including nouns, verbs, adjectives, or more abstract functional tokens pertinent to the context of the instruction. As illustrated in Figure 2, consider the meta template, “<verb> me <answer> to the question <related> the <image>: {question}”, where each placeholder is associated with a predefined set of positional synonyms, such as <verb> corresponds to three different candidates: “give”, “provide”, and “offer”. When generating templates, each placeholder is randomly assigned a candidate, allowing for diverse instruction templates to be produced. For example, one possible generated template is, “give me a response to the question concerning the provided image: {question}”
@@ 2.2 Diverse Template Sampling
Sentence pattern tree. We build a sentence pattern tree to systematically organize our instruction template space and diversely sample our templates. We use $T=(V,E)$ to denote the sentence pattern tree, where $V$ is the set of sentence patterns and $E$ is the edge between related sentence patterns. $T$ consists of four levels, ranging from coarse-grained to fine-grained, according to the taxonomy of sentence patterns. Level 1 represents the highest level of a sentence pattern, including declarative and imperative sentences. Level 2 decomposes Level 1 into simple, complex, and compound sentences. Level 3 further breaks Level 2 into subject-predicate, subject-predicate-object, subject-subject, noun clause, gerund clause, and linking clauses. Leaves in the final level represent the meta templates belonging to the above parent nodes. We then perform weighted sampling on Level 4 according to vertice features in Level 1 to Level 3. We construct two sentence pattern trees, one consists of 24 meta templates for visual questions and another consists of 14 meta templates for choices, yielding an extensive template space encompassing 15K visual question templates and 249K choice-related templates. We present the details of our sentence pattern trees with diverse meta templates in Appendix 7.1.

### v3

# Investigating the Scaling Effect of Instruction Templates for Training Multimodal Language Model

@@ Abstract
Current multimodal language model (MLM) training approaches overlook the influence of instruction templates. Previous research deals with this problem by leveraging hand-crafted or model-generated templates, failing to investigate the scaling effect of instruction templates on MLM training. In this work, we propose a programmatic instruction template generator capable of producing over 15K unique instruction templates by filling randomly sampled positional synonyms into weighted sampled meta templates, enabling us to comprehensively explore MLM’s performance across various template scales in the training process. Our investigation into scaling instruction templates for MLM training demonstrates that MLM capabilities do not consistently improve with increasing template scale. Instead, optimal performance is achieved at a medium template scale. Models trained with data augmented at the optimal template scale achieve performance gains of up to 10% over those trained on the original data and achieve the best overall performance compared with the similar-scale MLMs tuned on at most 75 times the scale of our augmented dataset. The code will be publicly available at https://github.com/shijian2001/TemplateScaling.
@@ 1 Introduction
Multimodal Language Models (MLMs) have revolutionized vision-language learning by performing visual instruction tuning on diverse, high-quality multimodal instruction data (Liu et al., 2024c; Zhu et al., 2023; Li et al., 2024; Zhang et al., 2024b). However, previous studies (Zhang et al., 2024a; Liu et al., 2024e; Sclar et al., 2023) reveal a critical limitation: MLMs exhibit substantial performance variability across different instruction templates. For instance, a succinct instruction and a detailed instruction can yield performance gaps exceeding 40% (Zhang et al., 2024a). This pronounced sensitivity to instruction templates compromises the reliability of MLM evaluation and diminishes the practical utility of MLMs in downstream applications.
Recent studies have empirically demonstrated that incorporating multiple instruction templates during MLM’s training process improves model performance and reduces instruction sensitivity (Zhang et al., 2024b; Razavi et al., 2025; Sanh et al., 2021). However, existing approaches primarily depend on either human-designed or model-generated small-scale templates, which suffer from limitations such as high costs, inherent design biases, and limited diversity in instruction formulations. Considering the success of scaling up training data significantly improves model’s performance (Fan et al., 2024; Kaplan et al., 2020) and the fact that multi-template training can improve MLM, this raises a critical question: How many instruction templates should be used during training to optimize MLM performance?
Figure 1: Training with the optimal template scale significantly improves MLM’s performance and reduces the performance variance. LLaVA-1.5-7B trained with 5K templates and LLaVA-1.5-13B trained with 100 templates achieve the highest average performance and the lowest performance variance among similar-scale MLMs on the SeedBench dataset, evaluated across 25 held-out instruction templates that are not included in training.
@@
To systematically investigate the scaling effect of instruction templates for MLM training, we propose a programmatic instruction template generator that leverages diverse meta templates to produce semantically equivalent instruction templates automatically and scalably. Our template generator can construct diverse instruction templates by random sampling from carefully curated word and phrase spaces to populate predefined placeholders, enabling the efficient generation of semantically consistent yet diverse instruction templates at scale. Our method can produce an extensive template space comprising 15K visual instruction templates. To ensure the diversity of sampled instruction templates from our template generator, we use a sentence-pattern tree organizational framework based on grammatical structures complemented by an efficient diverse sampling algorithm. This programmatic approach ensures the generation of instruction templates that maximize diversity across multiple dimensions, including grammatical construction, lexical choice, and symbolic representation.
Leveraging our programmatic template generator, we finetune two widely-used MLMs (LLaVA-1.5-7B and LLaVA-1.5-13B) (Liu et al., 2024a) and conduct a series of experiments by performing visual instruction tuning on the same dataset while varying the scale of instruction templates (from 10 to 15K). Our study reveals that the performance of MLMs does not consistently improve with the increasing scale of instruction templates. Instead, MLMs achieve the best general capabilities at a medium template scale, which varies with the model’s parameter size. We find LLaVA-1.5-7B’s performance peaks at 5K templates and LLaVA-1.5-13B peaks at 100 templates. We further compare our models trained under the optimal template scale with other MLMs fine-tuned on a significantly larger scale—up to 75.19 times the size of our instruction tuning datasets. Evaluation across five benchmarks reveals that our tuned models achieve the best overall performance (We showcase the comparison results on the SeedBench (Li et al., 2023b) dataset in Figure 1), thereby demonstrating the capacity of training with appropriate template scale to enhance MLMs in a data-efficient and cost-effective manner. Additionally, our analysis reveals that, compared to the original model, fine-tuning with the optimal template scale results in a substantial reduction in performance variance across various out-of-domain instruction templates. Our approach not only confirms the practical utility of the scaling effect of instruction templates but also provides promising insights into efficient strategies for improving MLMs. We summarize our main contributions as follows.
@@
(1) We introduce a novel programmatic instruction template generator that enables fast and scalable generation of diverse, semantically equivalent instruction templates.
(2) We comprehensively investigate the scaling effect of instruction templates for MLM training, demonstrating that MLM capabilities do not monotonically improve with increasing template scale and instead peak at a medium template scale.
@@
(3) We propose a simple yet effective approach to enhance visual instruction tuning by augmenting the original instruction tuning dataset with the optimal scale of templates we investigated. Our extensive experiments demonstrate its effectiveness.
@@ 2 Programmatically Scaling Instruction Templates
To investigate the scaling effect of instruction templates in MLM’s visual instruction tuning, we propose a programmatic instruction template generator. Our template generator can efficiently produce diverse, grammatically correct, and semantically consistent instruction templates. Specifically, we generate instruction templates by programmatically filling the pre-defined placeholders in a meta template with randomly sampled positional synonyms (phrases), ensuring flexibility and diversity while keeping the original meaning (Sec. 2.1). We organize our meta templates in a sentence pattern tree, along with a diverse template sampling algorithm to ensure the sampling probability across all instruction templates is uniformly distributed (Sec. 2.2).
@@ 2.1 Meta Templates
We design meta template $p_{i},i\in\{1,...,N\}$ as a formal blueprint for constructing instruction templates, consisting of a sequence of fixed string segments interspersed with placeholder $\langle h^{(i)}_{j}\rangle,j\in\{1,...,M_{i}\}$, where $M_{i}$ is the number of placeholders. We associate each placeholder $\langle h^{(i)}_{j}\rangle$ with a predefined set of synonyms (phrases) $s_{j}^{(i)}$. We design $s_{j}^{(i)}$ according to the semantic position of $\langle h^{(i)}_{j}\rangle$, including nouns, verbs, adjectives, or more abstract functional tokens pertinent to the context of the instruction. The potential template variations $\mathcal{T}(p_{i})$ grow combinatorially as $\mathcal{T}(p_{i})=\prod_{j=1}^{M_{i}}|s_{j}^{(i)}|$, where $|s_{j}^{(i)}|$ is the size of each synonym set. Fixed strings establish the foundational sentence structure, ensuring grammatical correctness and semantic coherence, while placeholders introduce flexibility and diversity, enabling the rapid generation of varied, high-quality instruction templates. We present the example of instruction template generation through a meta template in Figure 2. To ensure the diversity of generated visual instruction templates, we design 24 meta templates, yielding a template space capable of producing 15K distinct instruction templates.
@@ 2.2 Diverse Template Sampling
Sentence pattern tree. We build a sentence pattern tree to systematically organize our meta templates. We use $T=(V,E)$ to denote the sentence pattern tree, where $V$ is the set of sentence patterns and $E$ is the edge between related sentence patterns. $T$ consists of four levels, ranging from coarse-grained to fine-grained, according to the taxonomy of sentence patterns. We use level 1 to represent the highest level of a sentence pattern, including declarative and imperative sentences. Level 2 decomposes Level 1 into simple, complex, and compound sentences. Level 3 further breaks Level 2 into subject-predicate, subject-predicate-object, subject-subject, noun clause, gerund clause, and linking clauses. Leaves in the final Level 4 represent the meta templates belonging to the above parent nodes. Building on the tree framework, we can perform weighted sampling on Level 4 according to vertex features from Level 1 to Level 3. We present the details of our sentence pattern trees with diverse meta templates in Appendix A.